OpenStudy (anonymous):

On the moon the acceleration due to gravity is 1.6 metres per second squared (approximately 1 6 th of the value on earth). Standing on top of a ladder, 5 metres up, the astronaut throws a ball up vertically into the air with velocity 2m/s. How long does it take to reach the ground? How long would it take to reach the ground if the same experiment were done on earth?

OpenStudy (anonymous):

look in google

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

THAT DOESN'T HELP!

OpenStudy (anonymous):

I have an answer, but I am not entirely sure if it's correct or not. I'll write it down anyway.

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

\[\Delta x=v_i t+{1 \over 2}at^2\], where delta x=-5 (the change in the position), v_i=2m/s and a=-1.6m/s^2. Substituting them in the equation gives: \[-5=2t+{1 \over 2}(-1.6)t^2 \implies 0.8t^2-2t-5=0\]

OpenStudy (anonymous):

Solve the quadratic equation for t, you get one positive value for t. That's t=4.05 sec

OpenStudy (anonymous):

Do you have the answer, or choices?

OpenStudy (anonymous):

i only got to having to use the quadratic equation, i didn't realise i had to use the quadratic equation

OpenStudy (anonymous):

You can just do it in your calculator.

OpenStudy (anonymous):

yeah i just had to recognise that cause i think i'm meant to show working

OpenStudy (anonymous):

the quadratic formula is your personal picture yo :P

OpenStudy (anonymous):

yeah, cause i m having maximum trouble with it :)

OpenStudy (anonymous):

LOL!